Shift report 20000627 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jun 27 2000 - 08:16:32 EDT

  • Next message: http account: "Shift report 20000627 08:00-16:00"

    ------------------------------------------------------------
    Date: Tue Jun 27 08:16:31 2000, Shift: 00:00-08:00
    Supervisor: KH, with: YB
    Log book#2 pages 83 - 85
    
    Shift Summary:
    ------------------------------------------------------------
    Machine was down until 4:45
    They had miscelaneous problems.  In several instances
    they had things almost ready to go, but then something
    would dump.
    
    MCR called at 4:45 saying the beam was cogged at 70 GeV/u
    Started a run and got collisions (estimate 4000 from 
    5:00 - 8:00)
    Started run 1554 for the rest of the shift.
    Data rate was low (about 6-10/sec) and they were working
    first on steering then on a venier scan at STAR.  I don't
    know if that hurt our data rate or not.
    
    TPC2 Anode wires tripped several times
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jun 27 2000 - 08:20:46 EDT